5. A 10-m long plank (of negligible mass) is supported at each end by vertical cables. A person of unknown weight sits on the Can you explain the approach to solving a problem like this? Thank you!
Answer #1

SAY 10m

Let x be the distance of the person from the left end.

First, find the weight of the person

Consider all vertical forces

Fy=0

T +TR - W = 0

T +TR = 0

300N + 200N = W

W = 500M

==============

Take moment about the left end

Σ M, = 0

TR * 10m - W *I=0

TR * 10m = W *.

TR * 10m

\frac{200N*10m}{500N}=x

{\color{Red} x=4m}

ANSWER: C. 4.0 m

=================